OpenStudy (anonymous):

THE FACTORIZATION OF 64X^2+48X+9

OpenStudy (anonymous):

False. (64x + 3)(x + 3) = 64x^2 + 195x + 9

OpenStudy (anonymous):

DAFAQ SO CAN YOU HELP ME

OpenStudy (anonymous):

64x^2 + 48x + 9 = (8 x+3) (8 x+3) = (8 x+3)^2

OpenStudy (anonymous):

OH U BREAKDOWN THE 64 HUH

OpenStudy (anonymous):

yes! √64 = 8

OpenStudy (anonymous):

THANKS DAWGG